# Training towards significance with the decorrelated event classifier transformer neural network

### v1

@@ 4.4 Machine learning technique evaluation metrics
The significance is calculated using the method described in section 3.3, where the sample is divided into 8 machine learning technique bins with equal number of signal events in each bin.
The AUC curve is calculated using the scikit-learn Sklearn package with negative weights set to zero and evaluated in the Higgs boson mass window from 120 to 130 GeV.
@@
4.5 Experiment and results
A repeated random subsampling validation procedure RandomSubsample is used to evaluate the machine learning techniques, where 3 trials are done. The number of trials is limited due to the long training time of the neural networks. The validation procedure follows the steps below,
@@ 4.5 Experiment and results
  • The deep feed-forward network and event classifier transformer trained with BCE loss show the highest AUC and significance but have higher background background correlations.
@@
The machine learning technique metrics are shown for one of the subsampling trials. The correlation between the $m_{\ell\ell\gamma}$ and network output is shown in figure 5 for the deep feed-forward network trained with BCE loss and the event classifier transformer trained with the extreme + DisCo loss. The network output distribution is shown in figure 6.
Technique Loss Signi. AUC (%) Bkg. Corr. Best epoch
($10^{-3}$)
@@
Table 1: Average evaluation metrics for machine learning techniques using the random subsampling evaluation procedure with 3 trials. Random is a random classifier. FNN is feed-forward network. DFNN is deep feed-forward network. ETN is event classifier transformer network. N.A. is not available. Ext is extreme loss. Signi. is expected significance. AUC is area under the curve of the receiver operating characteristic curve. Bkg. Corr. is the correlation of $m_{\ell\ell\gamma}$ with the machine learning technique output for the background. Best epoch is the epoch that had the highest significance on the validation data set. Bold numbers indicate the best values over the machine learning techniques.
Figure 5: $m_{\ell\ell\gamma}$ distribution of the background, where each histogram is a bin in the machine learning technique output distribution. Each bin has an equal number of signal events. Lower signal percentile (sig. p) values correspond to higher network output values. (Left) Deep feed-forward network trained with BCE loss. (Right) Event classifier transformer network trained with extreme + DisCo loss. Correlation represents the magnitude of difference in the shapes between the machine learning bins. A lower correlation can be observed with the network trained with DisCo loss.
Figure 6: (Left) Deep feed-forward network trained with BCE loss. (Right) Event classifier transformer network trained with the extreme + DisCo loss. (Top) Network output distribution on the training data set and validation data set. (Bottom) Overtraining of the network evaluated by comparing the network output distribution between the training and validation data set for the signal and background. The comparison is done with the $\chi^{2}$ test Chi2Test implemented in ROOT ROOT1 , where residuals from the test are shown. For both networks, the bump in the middle of the signal distribution is due to the networks making different output distributions depending on $p_{T}^{\ell\ell\gamma}/m_{\ell\ell\gamma}$. An effect from the DisCo term can be seen in the right plot, where the minimum network output value is shifted upwards.
@@
During training of the feed-forward network with BCE loss, the BCE loss value and the significance can have different trends over the training epochs, which is shown in figure 7. This discrepancy demonstrates that when selecting the best model between the epochs, the expected significance should be used instead of the loss.
@@
5 Related work
The Particle Transformer ParticleTransformer is a transformer-based neural network targeted towards jet-flavor tagging. The structure is similar to the transformer Transformer , where each particle in a jet is considered as a token. Additionally, variables calculated using the features of a pair of particles are passed through a feed-forward network and added to the attention scores of the MHA. For the final prediction, class tokens are passed as a query to an MHA. The event classifier transformer in this work has a similar structure where the main difference is that each event feature is embedded to a separate token using separate feed-forward networks and that the neural network output is used to bin a bump hunt analysis.
There has been work to develop a loss function targeted towards significance, where a loss based on the inverse of the significance SignificanceLoss has been studied. However when a network is trained with this loss, the neural networks outputs are clustered around 0 or 1. This behavior makes the loss unusable for binning a bump hunt analysis. The proposed extreme loss can enhance the significance, while not having the network outputs clustered around 0 or 1. It can also reduces the number of epochs that are needed to reach the optimal performance of the network.
DisCo DisCo has been used for jet flavor tagging and for the ABCD analysis technique DiscoABCD , where it has been shown to be effective in decorrelating a targeted feature of the jet with the neural networks output and decorrelating two neural network outputs used for the ABCD method. In this work, DisCo is used to decorrelate the neural network output with the reconstructed mass to bin a bump hunting analysis.
6 Summary and conclusions

### v2

@@ IV.4 Machine learning technique evaluation metrics
The significance is calculated using the method described in section III.3, where the sample is divided into 8 machine learning technique bins with equal number of signal events in each bin.
The AUC curve is calculated using the scikit-learn [36] package with negative weights set to zero and evaluated in the Higgs boson mass window from 120 to 130 GeV.
@@
IV.5 Experiment and results
A repeated random subsampling validation procedure [37] is used to evaluate the machine learning techniques, where 3 trials are done. The number of trials is limited due to the long training time of the neural networks. The validation procedure follows the steps below,
@@ IV.5 Experiment and results
  • The deep feed-forward network and event classifier transformer trained with BCE loss show the highest AUC and significance but have higher background correlations.
@@
Table 1: Average evaluation metrics for machine learning techniques using the random subsampling evaluation procedure with 3 trials. Random is a random classifier. FNN is feed-forward network. DFNN is deep feed-forward network. ETN is event classifier transformer network. N.A. is not available. Ext is extreme loss. Signi. is expected significance. AUC is area under the curve of the receiver operating characteristic curve. Bkg. Corr. is the correlation of $m_{\ell\ell\gamma}$ with the machine learning technique output for the background calculated with the method described in section IV.4. Best epoch is the epoch that had the highest significance on the validation data set. Bold numbers indicate the best values over the machine learning techniques.
Technique Loss Signi. AUC (%) Bkg. Corr. Best
($10^{-3}$) epoch
@@
There is an upper limit on performance for a classifier due to the similarities of the signal and background, where machine learning classifiers could be optimized in approaching the upper limit in a certain phase spaces. Therefore the goal for the machine learning classifier is to approach the upper limit in the phase space that is most relevant for the problem. The event classifier transformer tends to be able to approach the upper limit on the significance metric better than the other machine learning techniques used in this paper. However with more hypertraining on the other machine learning techniques or by using different machine learning techniques, it could be possible to get closer to the upper limit, where this is left for future studies.
The machine learning technique metrics are shown for one of the subsampling trials. The correlation between the $m_{\ell\ell\gamma}$ and network output is shown in Fig. 5 for the deep feed-forward network trained with BCE loss and the event classifier transformer trained with the extreme + DisCo loss. The network output distribution is shown in Fig. 6, where in the figure overtraining of the network is measured with the $\chi^{2}$ test [38] implemented in ROOT [39], by comparing the network output distributions on the training sample and validation sample. The residuals in the plot are the normalized differences between the output distributions defined in [38]. For both networks, the bump in the middle of the signal distribution is related with the networks making different output distributions depending on $p_{T}^{\ell\ell\gamma}/m_{\ell\ell\gamma}$. An effect from the DisCo term can be seen, where the minimum network output value is shifted upwards.
Figure 5: $m_{\ell\ell\gamma}$ distribution of the background, where each histogram is a bin in the machine learning technique output distribution. Each bin has an equal number of signal events. Lower signal percentile (sig. p) values correspond to higher network output values. (Above) Deep feed-forward network trained with BCE loss. (Below) Event classifier transformer network trained with extreme + DisCo loss. Correlation represents the magnitude of difference in the shapes between the machine learning bins. A lower correlation can be observed with the network trained with DisCo loss.
Figure 6: (Left) Deep feed-forward network trained with BCE loss. (Right) Event classifier transformer network trained with the extreme + DisCo loss. (Top) Network output distribution on the training data set and validation data set. (Bottom) Overtraining of the network evaluated by comparing the network output distribution between the training and validation data set for the signal and background using a $\chi^{2}$ test, where residuals are the normalized differences between the output distributions.
@@
During training of the feed-forward network with BCE loss, the BCE loss value and the significance can have different trends over the training epochs, which is shown in Fig. 7. This discrepancy demonstrates that when selecting the best model between the epochs, the expected significance should be used instead of the loss.
@@
V Related work
The Particle Transformer [8] is a transformer-based neural network targeted towards jet-flavor tagging. The structure is similar to the transformer [7], where each particle in a jet is considered as a token. Additionally, variables calculated using the features of a pair of particles are passed through a feed-forward network and added to the attention scores of the MHA. For the final prediction, class tokens are passed as a query to an MHA. The event classifier transformer in this work has a similar structure where the main difference is that each event feature is embedded to a separate token using separate feed-forward networks and that the neural network output is used to bin a bump hunt analysis.
There has been work to develop a loss function targeted towards significance, where a loss based on the inverse of the significance [40] has been studied. However when a network is trained with this loss, the neural networks outputs are clustered around 0 or 1. This behavior makes the loss unusable for binning a bump hunt analysis. The proposed extreme loss can enhance the significance, while not having the network outputs clustered around 0 or 1. It can also reduces the number of epochs that are needed to reach the optimal performance of the network.
DisCo [15] has been used for jet flavor tagging and for the ABCD analysis technique [41], where it has been shown to be effective in decorrelating a targeted feature of the jet with the neural networks output and decorrelating two neural network outputs used for the ABCD method. In this work, DisCo is used to decorrelate the neural network output with the reconstructed mass to bin a bump hunting analysis.
VI Summary and conclusions